OpenStudy (anonymous):

Divide and check: (3x^2 − 7x + 1) ÷ (3x − 1)

OpenStudy (anonymous):

if you are lucky the numerator factors. try \(3x^2-7x+1)=(3x-1)\times (\text{something})\)

OpenStudy (anonymous):

oops i guess not you have to actually divide !

OpenStudy (anonymous):

you are going to get \[(x-2)-\frac{1}{3x-1}\]